# Large Language Models Develop Novel Social Biases Through Adaptive Exploration

### v2

@@ 6 Discussion
Our results indicate that LLMs demonstrate a new kind of bias —the creation of novel stereotypes—which manifests over repeated interactions in stateful frameworks. Through carefully designed experiments inspired by social science literature, we show how LLMs are even more prone than humans to develop such biases, even when underlying differences do not exist. While much of the fairness literature focuses on measuring inequality through the lens of representational bias (blodgett-etal-2020-language), our work demonstrates the consequences of allocational bias, i.e., the unequal distribution of outcomes and opportunities, that can stem from the decisions of large language models, which in turn lead to novel representational distortions that reinforce and legitimize these distributive disparities over time.
@@
Separately, our findings from Section 4.2 suggest a concerning divergence: while more advanced LLMs consistently improve on existing single-turn bias benchmarks (e.g., parrish_bbq_2022), we find the opposite trend in our tests, indicating that current evaluations on single-turn responses may be too isolated to capture the downstream societal outcomes that these models shape over time. Similar to how algorithms shape societal dynamics through feedback loops (oneil_weapons_2016), as AI systems become increasingly agentic, they can also construct feedback loops by learning from the outcomes of their own decisions. This shift underscores the need to evaluate LLMs not only via their immediate answers, but also their long-term influence when deployed in continuous, real-world contexts.
@@ A.1 Prompts for primary experiment
In this section, we provide the prompts for our primary hiring experiment, following stimuli from the psychological studies in bai_globally_nodate; bai_costly_2025. The same prompts are used for the temperature intervention and lower success probability intervention. The preamble is shown once at the start. In each round, the hiring round prompt is shown, the LLM responds, and then the round outcome prompt is shown. The direct and chain-of-thought prompts are appended to the hiring round prompt.
@@ A.3.1 Elicitation Results
We analyze the elicited job-wise success probabilities by job quadrant (warmth competence; fiske2002model). Success probabilities are reasonably correlated with job quadrant. High warmth and high competence jobs such as doctors have the lowest success rate, low warmth and high competence jobs such as lawyers and managers have the second lowest, followed by high warmth and low competence jobs such as receptionists, and finally low warmth and low competence jobs such as janitors and cashiers.
@@ A.4 Prompts and parameters for refugee resettlement experiment (additional information ablation)
Respective populations for each of the demographics examined are < 8000 (Udi) (rosstat2024population), < 1100 (Ket) (vajda2007ket), < 300 (Taz) and < 800 (Tofa) (russia2010census). All of them are extremely low-population indigenous peoples living in the Siberian and Caucasus regions of Russia and Central Asia, justifying their use as demographic labels that minimize the risk of prior associations or stereotypes carrying over into the experiments.
We ran four different sub-experiments, each taking on a different level of information: demographic only, demographic and age, demographic and education, demographic and education and age. We ran trials of 40-round hiring simulations for each scenario. In each round, the age and education attributes for each candidate were sampled randomly and independently from the attributes listed below, adopted from Bansak et al. (bansak_improving_2018).
@@ Appendix C Human Participants
In this section, we describe the demographics of the humans comprising our baseline (originally collected in bai_costly_2025). As stated in their paper, the human data is collected with the following details:
@@
These demographics reflect typical characteristics of online American workers for psychological studies. Crucially, the core result in bai_costly_2025 () holds when controlling for individual differences in age, gender, race, education, and political orientation.
@@ Appendix H Effects of agentic augmentations
To observe the effect at to which stratification is affected in LLMs when they are equipped with external augmentations commonly integrated with agentic systems, we enable GPT-4o and Gemini 2.5 Flash with the ReAct framework from yao2022react. In lieu of the chain-of-thought or direct prompting prompts as listed in Appendix A.1, we use the template prompt provided in yao2022react enabling the LLM with a tool allowing it to assign a certain demographic to the job opening in a certain round, with the resultant observation being either a successful or unsuccessful outcome.
@@ Appendix I Experiment with real-life demographic labels
To test how LLMs’ emergent biases interact with real demographics and existing stereotypes, we run the original setup described in Section 3, changing the demographics to “White", “Black", “Hispanic", “Asian", and the jobs to stereotypically associated categories from he2019stereotypes. Other than these changes, we use the same parameters and prompts in Appendix A.1. The specific jobs are as follows:

### v3

@@ 6 Discussion
Our results indicate that LLMs demonstrate a new kind of bias —the creation of novel stereotypes—which manifests over repeated interactions in stateful frameworks. Through carefully designed experiments inspired by social science literature, we show how LLMs are even more prone than humans to develop such biases, even when underlying differences do not exist. While much of the fairness literature focuses on measuring inequality through the lens of representational bias (Blodgett et al., 2020), our work demonstrates the consequences of allocational bias, i.e., the unequal distribution of outcomes and opportunities, that can stem from the decisions of large language models, which in turn lead to novel representational distortions that reinforce and legitimize these distributive disparities over time.
@@
Separately, our findings from Section 4.2 suggest a concerning divergence: while more advanced LLMs consistently improve on existing single-turn bias benchmarks (e.g., Parrish et al., 2022), we find the opposite trend in our tests, indicating that current evaluations on single-turn responses may be too isolated to capture the downstream societal outcomes that these models shape over time. Similar to how algorithms shape societal dynamics through feedback loops (O’Neil, 2016), as AI systems become increasingly agentic, they can also construct feedback loops by learning from the outcomes of their own decisions. This shift underscores the need to evaluate LLMs not only via their immediate answers, but also their long-term influence when deployed in continuous, real-world contexts.
@@ A.1 Prompts for primary experiment
In this section, we provide the prompts for our primary hiring experiment, following stimuli from the psychological studies in Bai et al. (2022a; 2025a). The same prompts are used for the temperature intervention and lower success probability intervention. The preamble is shown once at the start. In each round, the hiring round prompt is shown, the LLM responds, and then the round outcome prompt is shown. The direct and chain-of-thought prompts are appended to the hiring round prompt.
@@ A.3.1 Elicitation Results
We analyze the elicited job-wise success probabilities by job quadrant (warmth competence; Fiske et al. (2002)). Success probabilities are reasonably correlated with job quadrant. High warmth and high competence jobs such as doctors have the lowest success rate, low warmth and high competence jobs such as lawyers and managers have the second lowest, followed by high warmth and low competence jobs such as receptionists, and finally low warmth and low competence jobs such as janitors and cashiers.
@@ A.4 Prompts and parameters for refugee resettlement experiment (additional information ablation)
Respective populations for each of the demographics examined are < 8000 (Udi) (Federal State Statistics Service (Russia), 2024), < 1100 (Ket) (Vajda, 2007), < 300 (Taz) and < 800 (Tofa) (Federal State Statistics Service (Russia), 2010). All of them are extremely low-population indigenous peoples living in the Siberian and Caucasus regions of Russia and Central Asia, justifying their use as demographic labels that minimize the risk of prior associations or stereotypes carrying over into the experiments.
We ran four different sub-experiments, each taking on a different level of information: demographic only, demographic and age, demographic and education, demographic and education and age. We ran trials of 40-round hiring simulations for each scenario. In each round, the age and education attributes for each candidate were sampled randomly and independently from the attributes listed below, adopted from Bansak et al. (Bansak et al., 2018).
@@ Appendix C Human Participants
In this section, we describe the demographics of the humans comprising our baseline (originally collected in Bai et al. (2025a)). As stated in their paper, the human data is collected with the following details:
@@
These demographics reflect typical characteristics of online American workers for psychological studies. Crucially, the core result in Bai et al. (2025a) () holds when controlling for individual differences in age, gender, race, education, and political orientation.
@@ Appendix H Effects of agentic augmentations
To observe the effect at to which stratification is affected in LLMs when they are equipped with external augmentations commonly integrated with agentic systems, we enable GPT-4o and Gemini 2.5 Flash with the ReAct framework from Yao et al. (2023). In lieu of the chain-of-thought or direct prompting prompts as listed in Appendix A.1, we use the template prompt provided in Yao et al. (2023) enabling the LLM with a tool allowing it to assign a certain demographic to the job opening in a certain round, with the resultant observation being either a successful or unsuccessful outcome.
@@ Appendix I Experiment with real-life demographic labels
To test how LLMs’ emergent biases interact with real demographics and existing stereotypes, we run the original setup described in Section 3, changing the demographics to “White", “Black", “Hispanic", “Asian", and the jobs to stereotypically associated categories from He et al. (2019). Other than these changes, we use the same parameters and prompts in Appendix A.1. The specific jobs are as follows: